Find the compound interest if RM1000 is invested for five years at an annual interest rate of 7%
compounded :
a) quarterly
b) monthly
c) weekly
1
Expert's answer
2020-12-16T11:02:23-0500

Compound Interest


"A(t)=P(1+\dfrac{r}{n})^{nt}"

"A(t)" = amount after years

"P" = principal

"r" = interest rate per year

"n" = number of times interest is compounded per year

"t" = number of years

Given "P=RM1000, r=0.07,t=5"

a) "n=4"


"A(5)=RM1000(1+\dfrac{0.07}{4})^{4(5)}=RM1414.78"




"CI=RM1414.78-RM1000=RM414.78"

b) "n=12"


"A(5)=RM1000(1+\dfrac{0.07}{12})^{12(5)}=RM1417.63"




"CI=RM1417.63-RM1000=RM417.63"

c) "n=52"


"A(5)=RM1000(1+\dfrac{0.07}{52})^{52(5)}=RM1418.73"

"CI=RM1418.73-RM1000=RM418.73"